OpenStudy (anonymous):

Multivariable Calculus question: Hi everyone! I'm having trouble with the concept of integration with respect to a parameter. Hope someone can help. This isn't a problem from the MIT course.

OpenStudy (anonymous):

Here is the problem in question:

hartnn (hartnn):

thats DUIS differentiation under integral sign heard of it ?

OpenStudy (anonymous):

No, I haven't. My professor did an example in class (he probably called it by a different name), but now I'm confused.

hartnn (hartnn):

can you find the derivative of 1/(a^2+x^2)^2 with respect to 'a' ?

OpenStudy (anonymous):

(-4a)/(a^2+x^2)^3 right?

hartnn (hartnn):

yeah

hartnn (hartnn):

sorry, i meant to ask you can you find the derivative of 1/(a^2+x^2) with respect to 'a' ? :P

OpenStudy (anonymous):

(-2a)/(a^2+x^2)^2 I think...

hartnn (hartnn):

i'll write the rule by that time, \(\large F(a)=\int \limits_a^bf(x,a)dx \\ \dfrac{d}{dx}F(a) =\int \limits_a^b \dfrac{\partial}{\partial a}f(x,a)dx\)

hartnn (hartnn):

and yes! thats correct :)

hartnn (hartnn):

so we have \(\large F(a)=\int \limits_a^bf(x,a)dx = \dfrac{\pi}{2a}\\ \dfrac{d}{dx}F(a) =\int \limits_a^b \dfrac{\partial}{\partial a}f(x,a)dx = \dfrac{d}{dx}( \dfrac{\pi}{2a}) = \int_0^\infty \dfrac{2adx}{(x^2+a^2)^2}\) is this confusing ?

hartnn (hartnn):

i missed a negative sign :P

OpenStudy (anonymous):

oh ouch...

hartnn (hartnn):

\(\large F(a)=\int \limits_a^bf(x,a)dx = \dfrac{\pi}{2a}\\ \dfrac{d}{dx}F(a) =\int \limits_a^b \dfrac{\partial}{\partial a}f(x,a)dx = \dfrac{d}{dx}( \dfrac{\pi}{2a}) = \int_0^\infty \dfrac{-2adx}{(x^2+a^2)^2}\) whichpart did u not get ?

OpenStudy (anonymous):

I'm just a bit confused...if we're manipulating everything wrt a, then why is everything in dx?

OpenStudy (anonymous):

d/dx(pi/2a) would just be zero...

hartnn (hartnn):

we are JUST INTEGRTATING w.r.t x and yeah, i made a typo there again! sorry :P

hartnn (hartnn):

\(\large F(a)=\int \limits_a^bf(x,a)dx = \dfrac{\pi}{2a}\\ \dfrac{d}{da}F(a) =\int \limits_a^b \dfrac{\partial}{\partial a}f(x,a)dx = \dfrac{d}{da}( \dfrac{\pi}{2a}) = \int_0^\infty \dfrac{-2adx}{(x^2+a^2)^2}\)

OpenStudy (anonymous):

so the answer would just be (pi/2)ln(a)?

hartnn (hartnn):

\(\large F(a)=\int \limits_a^bf(x,a)dx \\ \large \dfrac{\pi}{2a}= \int_0^\infty \dfrac{dx}{(x^2+a^2)}\) just to show the co-relation, and pi/2 ln a is the left side! do we need the right side integral ?

OpenStudy (anonymous):

oh shucks yeah...

hartnn (hartnn):

\(\large \int_0^\infty \dfrac{-2adx}{(x^2+a^2)^2} = \dfrac{\pi \ln a }{2}\) since we are integrating w.r.t x we can take out -2a \(\large -2a\int_0^\infty \dfrac{dx}{(x^2+a^2)^2} \dfrac{\pi \ln a }{2}\) i think now you can find the required integral

OpenStudy (anonymous):

oh duh...since we're integrating wrt x we can treat a like a constant.

OpenStudy (anonymous):

just rearrange a bit and you have the required answer. haha.

hartnn (hartnn):

yes!

OpenStudy (anonymous):

Thank you so much!!!

hartnn (hartnn):

you're welcome ^_^